The Campus Division of All-States Bank has assets of $2,800 million. During the past year, the division had profits of $275 million. All-States Bank has a cost of capital of 5 percent. Ignore taxes.

Required:

a. Compute the divisional ROI for the Campus Division. (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)

b. Compute the divisional RI for the Campus Division. (Enter your answer in dollars, not in millions.)
Answer #1

a. ROI = Profits / Assets = 275 / 2800 = 9.8%

b. RI = Profits - (Assets x Cost of capital) = 275 - 2800 x 5% = $135000000